OpenStudy (anonymous):

Which of the following is the simplified form of (x+4)/(2x+6)+ 3/(x^2-9)? (x-2)/(2(x-3)) (x-2)/(2(x+3)) (x+2)/(2(x-3))

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

\[\frac{ x+4 }{ 2\left( x+3 \right) }+\frac{ 3 }{ \left( x+3 \right)\left( x-3 \right) }=\frac{ \left( x+4 \right) \left( x-3 \right)+2*3}{ 2\left( x+3 \right)\left( x-3 \right) }\] \[=\frac{ x^2+4x-3x-12+6 }{ 2\left( x+3 \right)\left( x-3 \right) }=\frac{ x^2+x-6 }{ 2\left( x+3 \right) \left( x-3 \right)}\] \[=\frac{ x^2+3x-2x-6 }{ 2\left( x+3 \right) \left( x-3 \right)}=?\]